A toy train set has a circular track peace the inner radius of the piece is 6 cm one sector of the truck has an arc length of 33 cm on the inside and 55 cm on the outside. What is the width of the track?


1
Expert's answer
2021-01-06T17:27:36-0500

Let r1 and r2 be the radius of the inner circle and outer circle of the track respectively.

Let l1 and l2 be the length of the inner arc and outer arc of the sector of the track respectively.

Let "\theta" be the angle of the sector.


Given:

r1 = 6 cm

l1 = 33 cm

l2 = 55 cm


The length of an arc of a sector is given as:

l = ("\theta"/360) * 2"\pi" * r

where r is radius and "\theta" is sector angle.


Therefore,

l1 = ("\theta"/360) * 2"\pi" * r1

33 = ("\theta"/360) * 2"\pi" * 6

"\theta" = (33 * 360)/(2"\pi" * 6)

"\therefore" "\theta" = 315.12"\degree"


Similarly,

l2 = ("\theta"/360) * 2"\pi" * r2

55 = (315.12/360) * 2"\pi" * r2

r2 = (55 * 360)/(2"\pi" * 315.12)

"\therefore" r2 = 10.0002 cm

i.e., r2 "\approxeq" 10 cm


Width of track = r2 - r1 = 10 - 6

"\therefore" Width of track = 4 cm